OpenStudy (anonymous):

Find an equation for the hyperbola that satisfies the given conditions. Foci (0, ±12), vertices (0, ±6)

OpenStudy (dumbcow):

center at (0,0) c = 12 a = 6 a^2 +b^2 = c^2 36 + b^2 = 144 b^2 = 108 Equation: \[\frac{y^{2}}{36}-\frac{x^{2}}{108} = 1\]